OpenStudy (anonymous):

Find the solution to each system by the addition (elimination)method. 4s - 5t =18 6s + 10t = -8 I need the steps please.

OpenStudy (anonymous):

first, choose which one you want to eliminate, s or t? if you eliminate s, you'll get t if you eliminate t, you'll get s first, let's eliminate t to eliminate t, i need to multiplied the first equation by 2 so i have 10t on both equation 4s - 5t =18 multiplied by 2 is 8s - 10t =36 now i can do elimination 8s - 10t =36 6s + 10t = -8 -------------- + (do addition) 8s + 6s -10t + 10t = 36 -8 14s = 28 s = 28/14 s=2

OpenStudy (anonymous):

then plug s into one of the equations 4s - 5t =18 4(2)-5t=18 8-5t=18 -5t=10 t=-2

OpenStudy (anonymous):

feel free to ask me if there's something you dont understand

OpenStudy (anonymous):

the answer I found it says no solution is that right?

OpenStudy (anonymous):

no, the solution is s=2 and t=-2

OpenStudy (anonymous):

thanks.